OpenStudy (anonymous):

Did I do this right? q water = 24 ml x 4.18 J / (g ˚C) x (27.5 ˚C – 24 ˚C) q = 351 J water = -351 J unknown metal -351 J = 15.262 g x c x (27.5 ˚C – 100.3 ˚C) c = -351 / (15.262 x -72.8) c = 0.316 J (g ˚C) Mass of metal: 15.262 g Volume of water in the calorimeter: 24.0 mL Initial temperature of water in calorimeter: 25.2 °C Temperature of hot water and metal in hot water bath: 100.3 °C Final temperature reached in the calorimeter: 27.5 °C

OpenStudy (anonymous):

That is what it is asking and I can't find the right an answer that is close to it. Metal Specific Heat Capacity Nickel 0.440 Tin 0.210 Silver 0.237 Magnesium 1.020 Calcium 0.650 Mercury 0.140

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

Can you please post the question first?

OpenStudy (anonymous):

1. Calculate the energy change (q) of the surroundings (water) using the enthalpy equation 2. Using the formula q unknown metal = m × c × ΔT, calculate the specific heat of the metal. Use the data from your experiment for the unknown metal in your calculation. Show ALL your work. Given : Mass of metal: 15.262 g Volume of water in the calorimeter: 24.0 mL Initial temperature of water in calorimeter: 25.2 °C Temperature of hot water and metal in hot water bath: 100.3 °C Final temperature reached in the calorimeter: 27.5 °C

OpenStudy (anonymous):

1: q water = 24 ml x 4.18 J / (g ˚C) x (27.5 ˚C – 24 ˚C) q = 351 J water = -351 J unknown metal 2: -351 J = 15.262 g x c x (27.5 ˚C – 100.3 ˚C) c = -351 / (15.262 x -72.8) c = 0.316 J (g ˚C) But this answer doesn't overlap with the next question

OpenStudy (anonymous):

Which is this : 2. Using the specific heat capacity value that you determined in Part II of the lab, what is the most probable identity of the unknown metal that you examined? Please explain your reasoning in complete sentences. Based on your answer in question 2, determine the percent error of your calculated specific heat capacity value of your unknown metal using the specific heat capacities of known metals below. Metal Specific Heat Capacity Nickel 0.440 Tin 0.210 Silver 0.237 Magnesium 1.020 Calcium 0.650 Mercury 0.140

OpenStudy (anonymous):

@mathslover